# Sparse Multivariate Linear Regression with Strongly Associated Response Variables

### v2

@@ Appendix B Proof of Theorem 1
Under the conditions (A1) and (A3), assuming that $(\hat{\eta}^{2},\hat{\theta})$ are $\sqrt{n}$-consistent estimators of $(\eta_{*}^{2},\theta_{*})$, then, if $\lambda\sqrt{n}\rightarrow 0$ and $\lambda n^{(r+1)/2}\rightarrow\infty$, the following holds.
@@
where $\hat{B}$ is a solution to the optimization problem (14) for fixed $(\hat{\eta}^{2},\hat{\theta})$, and $D=(\Omega_{*}\otimes Z)_{S}$.
@@
The diagonals of $\Omega_{*}$ are $\frac{1+(q-2)\theta_{*}}{\eta^{2}_{*}(1-\theta_{*})(1+(q-1)\theta_{*})}=\frac{1}{\eta^{2}_{*}(1-\theta_{*})}(1-\frac{\theta_{*}}{1+(q-1)\theta_{*}})$ and the off-diagonals are $-\frac{\theta_{*}}{\eta^{2}_{*}(1-\theta_{*})(1+(q-1)\theta_{*})}$. Let $v_{*}=\frac{1}{\eta^{2}_{*}(1-\theta_{*})}$ and $w_{*}=\frac{\theta_{*}}{\eta^{2}_{*}(1-\theta_{*})(1+(q-1)\theta_{*})}$. We further denote $\hat{v}=\frac{1}{\hat{\eta}^{2}(1-\hat{\theta})}$ and $\hat{w}=\frac{\hat{\theta}}{\hat{\eta}^{2}(1-\hat{\theta})(1+(q-1)\hat{\theta})}$. Then,
$|\hat{v}-v_{*}|$ $=\left|\frac{\hat{\eta}^{2}(1-\hat{\theta})-\eta^{2}_{*}(1-\theta_{*})}{\eta^{2}_{*}(1-\theta_{*})\hat{\eta}^{2}(1-\hat{\theta})}\right|$
@@
We define a matrix, $V\in\mathbb{R}^{2\times 2}$, which characterizes the limiting covariance of $(\hat{\eta},\hat{\theta})$. This symmetric matrix has the following four elements:
@@
where $(\hat{\eta}^{2},\hat{\theta})$ is a solution to the optimization problem (14) for fixed $\hat{B}$. Furthermore, if $\theta_{*}\in(0,1)$, then
@@
It suffices to assume that true $B_{*}$ is known, since the natural following step which is the replacement of $B_{*}$ with $\hat{B}$ can be directly obtained by application of the derivation used in the proof of Theorem 2 in Lee and Liu [6]. Recall the reparametrization of $(\eta^{2},\theta)$ to $(\alpha,\gamma)$ via $\alpha=\eta^{2}(1-\theta)$, $\gamma=\eta^{2}(1+(q-1)\theta)$ by which we derived the closed form solutions to $\hat{\alpha}$ and $\hat{\gamma}$. The solutions were $\hat{\alpha}=\frac{qM_{1}^{*}-M_{2}^{*}}{q(q-1)}$, $\hat{\gamma}=\max\{\hat{\alpha},M_{2}^{*}/q\}$, where $M_{1}^{*}=\frac{1}{n}\|E\|_{F}^{2}$ and $M_{2}^{*}=\frac{1}{n}\|E1_{q}\|^{2}$ (see (6), (7)). We define $\hat{\delta}=\frac{M_{1}^{*}}{q}$. We first find the joint limiting distribution of $(\hat{\delta},\hat{\alpha})$. We have
@@
Then, by the same Delta method via $g(x,y)=(x,(x-y)/x)$, we have

### v3

@@ Appendix B Proof of Theorem 1
[…]
@@
[…]
@@
The diagonals of $\Omega_{*}$ are $\frac{1+(q-2)\theta_{*}}{\eta^{2}_{*}(1-\theta_{*})(1+(q-1)\theta_{*})}=\frac{1}{\eta^{2}_{*}(1-\theta_{*})}(1-\frac{\theta_{*}}{1+(q-1)\theta_{*}})$ and the off-diagonals are $\frac{\theta_{*}}{\eta^{2}_{*}(1-\theta_{*})(1+(q-1)\theta_{*})}$. Let $v_{*}=\frac{1}{\eta^{2}_{*}(1-\theta_{*})}$ and $w_{*}=\frac{\theta_{*}}{\eta^{2}_{*}(1-\theta_{*})(1+(q-1)\theta_{*})}$. We further denote $\hat{v}=\frac{1}{\hat{\eta}^{2}(1-\hat{\theta})}$ and $\hat{w}=\frac{\hat{\theta}}{\hat{\eta}^{2}(1-\hat{\theta})(1+(q-1)\hat{\theta})}$. Then,
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Then, by the same Delta method via g(x,y)=(x,(xy)/x)g(x,y)=(x,(x-y)/x), we have